OpenStudy (anonymous):

I need to know how to determine if the square root or 27/4 is a rational, irrational, or an interger?

OpenStudy (anonymous):

sqrt(27/4)= sqrt(3^2.3/2^2) = 3sqrt2/2 because sqrt 2 is irrational, therefore sqrt(27/4) irrational

OpenStudy (anonymous):

where did you get the (3^2.3/2^2)